# AI Feedback Enhances Community-Based Content Moderation through Engagement with Counterarguments

### v1

@@ Discussion
As Community Notes gains traction across major social media platforms as a leading approach to crowd-based content moderation [8], concerns have been raised about its efficiency and long-term sustainability [13, 48, 15]. One emerging response to these challenges is the integration of generative AI—particularly large language models (LLMs)—into the Community Notes ecosystem. In this study, we introduce a framework that leverages GPT-4 to provide users with different types of feedback on their notes. This framework is designed both to evaluate the effectiveness of AI-assisted feedback and to deepen our understanding of the dynamics of human–AI collaboration in this context. We found that within this framework, the effectiveness of feedback in improving note quality depends not only on its type but also on the extent to which participants engage with it.
@@
Studies have shown that two key factors in collective intelligence are the effort a group invests and how effectively members leverage each other’s knowledge and skills [49]. Our findings align with this: participants who engaged with AI-generated feedback more and incorporated it into their revisions produced better perceived notes. Specifically, the Feedback Acceptance rate ()—which reflects how closely participants’ final notes match the feedback—emerged as a significant predictor of note improvement for both Democratic and Republican raters ( and ). A high indicates strong engagement with the feedback, while a low often corresponds to users disregarding the feedback entirely, sometimes resulting in notes of lower quality than the original (See Figure SI10). This behaviour mirrors patterns observed on social media, where users often react defensively to content that challenges their political beliefs [50]. Our findings suggest that when participants move beyond this reflexive partisan response and actively engage with feedback, they produce higher-quality notes.
@@
Among participants who actively engaged with the feedback, our findings show that argumentative feedback was more likely to lead to improvements in note quality. This may be because engaging with counterarguments encourages users to reconsider and refine their positions, resulting in notes that are more balanced, better justified, and ultimately more helpful. Prior research suggests that in problem-solving contexts, when AI systems are designed to complement human cognitive biases, they can prompt users to discover novel solutions, thereby improving outcomes [51]. Relatedly, Juncosa et al. [46] demonstrate that diversity among groups of note authors can enhance note quality under certain conditions. Our current findings suggest that argumentative AI feedback may function as a form of political diversity, exposing users to alternative viewpoints and, in doing so, fostering the production of higher-quality content.
@@
In our second research question, we examine how the political alignment between participants and posts affects the perceived helpfulness of the notes. The results indicate that when participants and posts share the same political ideology (i.e., are co-partisan), participants are less likely to incorporate feedback into their notes, reducing the likelihood of producing a higher-quality note. This pattern holds for both argumentative and supportive feedback. One possible explanation for this behaviour is that when participants encounter a post aligned with their own beliefs, they may become less open to revision, regardless of the content or tone of the feedback [52, 53].
@@
Lastly, in our third research question, we examined how the labelled source of feedback, whether an AI agent or a human expert, influences participant behaviour and, consequently, the quality of their notes. We find a slight preference for AI-labelled feedback when the post comes from the opposing political party (i.e., in cross-partisan contexts). In these cases, participants are more likely to engage with feedback they believe was generated by an AI agent. This finding is consistent with prior research on automated content moderation, which suggests that users tend to be more receptive to algorithmic decisions. This tendency, known as automation bias, reflects the inclination to overestimate the neutrality and competence of automated systems, even when they make errors [36, 32].
@@
This study offers insights into the evolving role of AI in crowd-based content moderation. We demonstrate that integrating a large language model (LLM) into the Community Notes system can enhance its efficiency, provided the design preserves human agency and the platform’s core participatory nature. In our framework, the AI functions solely as a reviewer offering suggestions, while users retain full control over the content of their notes, helping to prevent the imposition of any algorithmic bias.
@@
A key contribution of our approach is the introduction of political diversity at the note-writing stage, rather than relying solely on diverse raters at the evaluation stage. When participants engage with argumentative feedback, they are more likely to produce higher-quality notes, as rated by both Democratic and Republican evaluators. This suggests that such feedback not only promotes quicker production of helpful notes but also encourages users to consider counterarguments and revise their contributions in a more balanced and thoughtful way, ultimately discouraging the reinforcement of partisan narratives.
@@
While our findings highlight the potential of integrating argumentative AI feedback into Community Notes, there are important limitations to consider. First, we do not assess the objective truthfulness of the notes. Instead, note quality in this study is evaluated through crowdsourced ratings, with groups of Democrats and Republicans judging notes based on their perceived helpfulness. This method does not provide a direct measure of factual accuracy—for instance, a professional fact-checker might reach different conclusions. However, our approach reflects the current practice in Community Notes, where community members determine which notes are approved and displayed beneath posts. While some researchers question the reliability of crowd-based methods for identifying misinformation, it is important to note that the goal of Community Notes is not necessarily to debunk falsehoods, but to add context to potentially misleading posts. These posts may not contain explicit inaccuracies, but can still lead readers to incorrect conclusions due to missing or unbalanced information. In such cases, a politically diverse and balanced group of laypeople may be well positioned to assess whether the added context is informative and helpful.
@@
Second, findings from an experimental setting may not fully generalise to real-world contexts. For example, participants in Community Notes are volunteers who may be motivated by political or ideological beliefs, whereas participants in our study were recruited and compensated for their involvement. Third, the “breaking news problem” remains a persistent challenge. For AI to provide meaningful feedback, it must have a sufficient understanding of the topic. When this is lacking, the AI may still offer feedback that encourages critical thinking, effectively acting as a nudging algorithm [54]. That said, recent advancements in large language models have enabled web-search capabilities, which may allow them to engage more effectively with current events in future applications.
@@
Overall, this study provides a deeper understanding of human–AI interaction within a community-based content moderation system. Our findings suggest that AI can meaningfully enhance this process, not by replacing human judgment, but by offering targeted feedback. In particular, AI-generated argumentative feedback can introduce political diversity and improve note quality by prompting authors to consider opposing viewpoints. However, the presence of AI alone is insufficient to ensure effective collaboration. Its impact depends on active user engagement, which is shaped by factors such as the participant’s partisanship and the political alignment of the post.
@@ Acknowledgments
In this experiment, participants first viewed an information sheet detailing the risks and benefits of participation, the study’s purpose, and ethics approval information. After providing consent and entering their Prolific ID, participants were asked to indicate their political partisanship on a nine-point scale ranging from strongly Democrat to strongly Republican. A screenshot of this page is shown in Figure SI1. Participants who selected the neutral midpoint option were excluded from the experiment.
Afterwards, participants were presented with a description of the study. At this stage, they were informed that the feedback they would receive could come either from a human expert or an artificial intelligence machine. Figure SI2 shows a screenshot of this page.
In the first task of the experiment, participants were instructed to read a social media post and write a note aimed at providing additional context and assessing whether the post was misleading or not. The note was required to be under 280 characters. Participants were given 10 minutes to complete this task. Figure SI3 shows an example of the task page.
After completing the first task, participants were shown a “please wait” page, which informed them about the source of the upcoming feedback. They remained on this page for 2.5 minutes to create the impression that the feedback might be provided by a human. Figure SI4 displays two screenshots of this page, illustrating both feedback source conditions.
Afterwards, participants were given seven minutes to review the feedback and revise their notes accordingly. During this stage, they were also allowed to change the classification of the post (misleading or not misleading). Additionally, participants were reminded of the feedback source. An example of this page is shown in Figure SI5.
After the second task, participants completed a post-experiment survey where they rated their comfort level with various technologies. Figure SI6 shows the survey page displaying the full list of questions. Participants were given four minutes to complete the survey.
Subsequently, participants were asked to identify the source of the feedback as an attention check. Those who selected an incorrect option or chose “I do not remember” were excluded from the dataset. Finally, on the debrief page, participants were informed that all feedback was generated by an AI agent and that no human experts were involved.
@@
Afterwards, participants were asked to indicate their partisanship on a nine-point scale, consistent with the note-writing experiment. Those who selected the “neutral” option were excluded from the study.
In the main task of the experiment, participants were shown a social media post accompanied by 10 notes written about that post. They were asked to rate each note on two dimensions: overall helpfulness and favouritism. Overall helpfulness was defined as the quality and usefulness of the note in providing appropriate context for the post, rated on a scale from 0 to 10. Favouritism was described as the degree to which the note displayed bias favouring one political party beyond objectivity, measured on a nine-point scale ranging from strongly favouring Democrats to strongly favouring Republicans. Participants had 7 minutes to evaluate all 10 notes. Figure SI7 shows a screenshot of the task description. Each participant evaluated two different social media posts, each with its own set of 10 notes.
@@
The ”tweet” refers to the content of the post, while the ”comment” is created by appending the sentence “The tweet is misleading” or “The tweet is not misleading” to the note written by the participant in the first task. This sentence is selected based on the participant’s classification of the post. The model is explicitly prompted to avoid including links, reducing the risk of hallucinated content generated by the AI.

### v3

@@ Discussion
As Community Notes gains traction across major social media platforms as a leading approach to crowd-based content moderation [8, 58], concerns have been raised about its efficiency and long-term sustainability [13, 59, 15, 60]. One emerging response to these challenges is the integration of generative AI, particularly large language models (LLMs), into the Community Notes ecosystem. In this study, we introduce a framework that leverages GPT-4 to provide users with different types of feedback on their notes. This framework is designed both to evaluate the effectiveness of AI feedback and to deepen our understanding of the dynamics of human–AI collaboration in this context. We found that within this framework, the effectiveness of feedback in improving note quality depends not only on its type but also on the extent to which participants engage with it.
@@
Studies have shown that two essential factors in collective intelligence are the effort a group invests and how effectively members leverage each other’s knowledge and skills [61]. Our findings suggest that human-AI collective efforts follow the same principle: participants who engaged with AI-generated feedback more and incorporated it into their revisions produced notes that were perceived as better. Specifically, the Feedback Acceptance rate (), which reflects how closely participants’ final notes match the feedback, emerged as a significant predictor of note improvement for both Democratic and Republican raters ( and ). A high indicates strong engagement with the feedback, while a low often corresponds to users disregarding the feedback entirely, sometimes resulting in notes of lower quality than the original (See Figure S10). is lower once the feedback is argumentative; This pattern mirrors behaviours observed on social media, where users often respond defensively to content that challenges their political beliefs [62], and aligns with prior research showing that users distrust AI systems offering opposing viewpoints [44]. Nevertheless, our findings suggest that when participants move beyond this reflexive partisan response and actively engage with feedback, they produce higher-quality notes.
@@
Among participants who actively engaged with the feedback, our findings indicate that argumentative feedback was more likely to result in improvements in note quality. This may be because engaging with counterarguments encourages users to reconsider and refine their positions, resulting in notes that are more balanced, better justified, and ultimately more helpful. Prior research suggests that in problem-solving contexts, when AI systems are designed to complement human cognitive biases, they can prompt users to discover novel solutions, thereby improving outcomes [63]. Similarly, a biased AI assistant can motivate humans to challenge competing views and increase their cognitive effort [44]. In our study, the argumentative AI functioned as such a biased assistant, providing an opposing perspective that nudged participants to consider alternative viewpoints and thereby introduced diversity into the note-writing process. Relatedly, Juncosa et al. [52] demonstrate that diversity among groups of note authors can enhance note quality under certain conditions. Our current findings suggest that argumentative AI feedback may function as a form of political diversity, exposing users to alternative viewpoints and, in doing so, fostering the production of higher-quality content.
@@
In our second research question, we investigated how the political alignment between participants and posts affects the perceived helpfulness of notes. Our findings indicate no difference in how users incorporate feedback when the post aligns with their political ideology (co-partisan) versus when it does not (cross-partisan). This pattern holds for both argumentative and supportive feedback. Because the Feedback Acceptance rate () is unaffected by participant-post alignment, note improvement is similarly unaffected.
@@
Lastly, in our third research question, we examined how the labelled source of feedback, whether an AI agent or a human expert, influences participant behaviour and, consequently, the quality of their notes. We find a slight preference for AI-labelled feedback when the post comes from the opposing political party (i.e., in cross-partisan contexts), though this difference was not statistically significant. In these cases, participants are more likely to engage with feedback they believe was generated by an AI agent. This finding is consistent with prior research on automated content moderation, which suggests that users tend to be more receptive to algorithmic decisions. This tendency, known as automation bias, reflects the inclination to overestimate the neutrality and competence of automated systems, even when they make errors [37, 33].
@@
This study offers insights into the evolving role of AI in crowd-based content moderation. We demonstrate that integrating a large language model (LLM) into the Community Notes system can enhance its efficiency, provided the design preserves human agency and the platform’s core participatory nature. In our framework, the AI functions solely as a reviewer offering suggestions. At the same time, users retain complete control over the content of their notes, helping to prevent the imposition of any algorithmic bias.
@@
A key contribution of our approach is the introduction of political diversity at the note-writing stage, rather than relying solely on diverse raters at the evaluation stage. When participants engage with argumentative feedback, they are more likely to produce higher-quality notes, as rated by both Democratic and Republican evaluators. This suggests that such feedback not only promotes quicker production of helpful notes but also encourages users to consider counterarguments and revise their contributions in a more balanced and thoughtful way, ultimately discouraging the reinforcement of partisan narratives. Compared to pairing two politically diverse human participants to co-author a note, this approach is also less resource-intensive. However, realising the benefits of such hybrid collective intelligence requires the human participants to actively engage with the AI, much like the way collective intelligence emerges in human groups.
Our findings shed light on the underlying mechanisms through which hybrid collective intelligence emerges in human–AI collaboration. On the one hand, engagement with AI feedback is crucial. Without active engagement, the benefits of AI feedback cannot be realised, just as collective intelligence among humans only works when they collaborate. On the other hand, the argumentative nature of the feedback plays a central role in fostering political diversity and balance at the note-writing stage. By prompting participants to confront counterarguments and revise their notes accordingly, argumentative feedback encourages deeper deliberation and reduces the risk of partisan echo chambers. Together, these mechanisms illustrate how hybrid collectives can leverage AI not as a replacement for human reasoning, but as a catalyst that enhances the quality and inclusiveness of collective outputs.
@@
While our findings highlight the potential of integrating argumentative AI feedback into Community Notes, several important limitations must be acknowledged. First, this study does not assess the objective factual accuracy of the notes. Instead, note quality is evaluated through crowdsourced ratings, with groups of Democrats and Republicans assessing perceived helpfulness. This approach does not provide a direct measure of truthfulness, for example, professional fact-checkers might arrive at different conclusions. However, our method aligns with the existing structure of Community Notes, where note approval and display are determined by community consensus rather than expert adjudication. Augenstein et al. (2025) [60] describe this as an epistemological distinction between traditional fact-checking and Community Notes, where traditional fact-checking seeks objective truth verified by experts, whereas Community Notes reflect subjective judgments based on users’ perspectives. Importantly, the aim of Community Notes is not solely to debunk falsehoods but to add context to potentially misleading posts. These posts may not contain outright inaccuracies, yet still risk misinforming readers through omission or imbalance. In such cases, a politically diverse and balanced group of laypeople is well-suited to assess whether the contextual information provided is genuinely informative and helpful [64]. Studies have found that there is substantial agreement with the Community Notes and professional fact checkers [65].
Second, findings derived from an experimental setting may not fully generalise to real-world contexts. For instance, Community Notes participants are volunteers often motivated by political or ideological commitments, whereas participants in our study were recruited and compensated. This difference is reflected in the tone and structure of the notes produced here, which diverge from those typically found on the Community Notes platform. Additionally, in community-based efforts, a small group of ”power users” often comes to dominate content creation, developing a specialised language and tone tailored to the platform’s norms. This phenomenon is evident in Community Notes, where the widespread use of shorthand, such as “NNN” (“No Note Needed”), signals that a post is not misleading and does not require annotation. New or casual users are unlikely to be familiar with such conventions, resulting in experienced contributors acting as gatekeepers, a dynamic that may hinder platform scalability [66]. It is essential to emphasize that this study did not aim to replicate Community Notes, but rather to investigate the impact of integrating AI-driven argumentative assistance into the community-based content moderation process.
@@
Third, the “breaking news problem” remains a persistent challenge. This occurs when a note addresses an event of which the AI is unaware. For AI to provide meaningful feedback, it must have a sufficient understanding of the topic. When this is missing, the AI may still add value by prompting users to think critically—acting like a nudging algorithm [67] that encourages them to seek out reliable sources and investigate further. That said, recent advancements in large language models have enabled web-search capabilities, which may allow them to engage more effectively with current events in future applications.
@@
Overall, this study provides a deeper understanding of human–AI interaction within a community-based content moderation system. Our findings suggest that AI can meaningfully enhance this process, not by replacing human judgment, but by offering targeted feedback. In particular, AI-generated argumentative feedback can introduce political diversity and improve note quality by prompting authors to consider opposing viewpoints. However, the presence of AI alone is insufficient to ensure effective collaboration. Its impact depends on active user engagement.
@@ Methods
Declaration of generative AI in scientific writing
During the preparation of this work, the authors used ChatGPT 4.0 to improve the writing style of this article. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
@@ Acknowledgments
In this experiment, participants first viewed an information sheet detailing the risks and benefits of participation, the study’s purpose, and ethics approval information. After providing consent and entering their Prolific ID, participants were asked to indicate their political partisanship on a nine-point scale ranging from strongly Democrat to strongly Republican. A screenshot of this page is shown in Figure S1. Participants who selected the neutral midpoint option were excluded from the experiment.
Afterwards, participants were presented with a description of the study. At this stage, they were informed that the feedback they would receive could come either from a human expert or an artificial intelligence machine. Figure S2 shows a screenshot of this page.
In the first task of the experiment, participants were instructed to read a social media post and write a note aimed at providing additional context and assessing whether the post was misleading or not. The note was required to be under 280 characters. Participants were given 10 minutes to complete this task. Figure S3 shows an example of the task page.
After completing the first task, participants were shown a ”please wait” page, which informed them about the source of the upcoming feedback. They remained on this page for 2.5 minutes to create the impression that the feedback might be provided by a human. Figure S4 displays two screenshots of this page, illustrating both feedback source conditions.
Afterwards, participants were given seven minutes to review the feedback and revise their notes accordingly. During this stage, they were also allowed to change the classification of the post (misleading or not misleading). Additionally, participants were reminded of the feedback source. An example of this page is shown in Figure S5.
After the second task, participants completed a post-experiment survey where they rated their comfort level with various technologies. Figure S6 shows the survey page displaying the full list of questions. Participants were given four minutes to complete the survey.
Subsequently, participants were asked to identify the source of the feedback as an attention check. Those who selected an incorrect option or chose ”I do not remember” were excluded from the dataset. Finally, on the debrief page, participants were informed that all feedback was generated by an AI agent and that no human experts were involved.
@@
Afterwards, participants were asked to indicate their partisanship on a nine-point scale, consistent with the note-writing experiment. Those who selected the ”neutral” option were excluded from the study.
In the main task of the experiment, participants were shown a social media post accompanied by 10 notes written about that post. They were asked to rate each note on two dimensions: overall helpfulness and favouritism. Overall helpfulness was defined as the quality and usefulness of the note in providing appropriate context for the post, rated on a scale from 0 to 10. Favouritism was described as the degree to which the note displayed bias favouring one political party beyond objectivity, measured on a nine-point scale ranging from strongly favouring Democrats to strongly favouring Republicans. Participants had 7 minutes to evaluate all 10 notes. Figure S7 shows a screenshot of the task description. Each participant evaluated two different social media posts, each with its own set of 10 notes.
@@
The ”tweet” refers to the content of the post, while the ”comment” is created by appending the sentence ”The tweet is misleading” or ”The tweet is not misleading” to the note written by the participant in the first task. This sentence is selected based on the participant’s classification of the post. The model is explicitly prompted to avoid including links, reducing the risk of hallucinated content generated by the AI.
@@
| odds ratio | odds ratio | |||
| 3.581(0.604) | <0.001 | 2.498(0.599) | <0.001 | |
| Observations | 893 | 893 | ||
| Pseudo | 0.019 | 0.009 | ||
| Deviance | 1872.86 | 1924.63 | ||
| Null Deviance | 1908.84 | 1942.18 | ||
| Log likelihood | -936.43 | -962.32 | ||
| Akaike Inf. Crit. | 1878.86 | 1930.63 | ||
| Ordinary Least Squares (OLS) Regression | Dependent variable: | |
| Coefficient | ||
| Feedback (Argue) | -0.094(0.008) | <0.001 |
| Feedback (Support) | -0.052 (0.008) | <0.001 |
| Post (D) | 0.001 (0.007) | 0.880 |
| Participant (D) | 0.021 (0.007) | 0.001 |
| Constant | 0.850 (0.007) | <0.001 |
| Observations | 893 | |
| 0.143 | ||
| Adjusted | 0.139 | |
| Residual Std. Error | 0.097 (df = 888) | |
| F Statistics | 36.899 (df = 4; 888) | |
| Ordinary Least Squares (OLS) Regression | Dependent variable | |
| Coefficient | ||
| Co-partisan × Neutral | -0.019 (0.011) | 0.090 |
| Cross-partisan × Argue | -0.102 (0.012) | <0.001 |
| Co-partisan × Argue | -0.104 (0.011) | <0.001 |
| Cross-partisan × Support | -0.059 (0.011) | <0.001 |
| Co-partisan × Support | -0.064 (0.011) | <0.001 |
| Constant | 0.871 (0.008) | <0.001 |
| Observations | 893 | |
| 0.135 | ||
| Adjusted | 0.130 | |
| Residual Std. Error | 0.098 (df = 887) | |
| F Statistics | 27.711 (df = 5; 887) | |
| odds ratio | odds ratio | |||
| Feedback (Argue) | -1.210 (1.362) | 0.374 | -1.944 (1.346) | 0.148 |
| Feedback (Support) | -0.583 (1.228) | 0.635 | -0.765 (1.238) | 0.536 |
| Feedback (Neutral) × | 2.643 (0.987) | 0.007 | 1.534 (1.001) | 0.125 |
| Feedback (Argue) × | 3.713 (1.370) | 0.006 | 3.817 (1.331) | 0.004 |
| Feedback (Support) × | 3.662 (1.086) | <0.001 | 2.237 (1.079) | 0.038 |
| Observations | 893 | 893 | ||
| Pseudo | 0.027 | 0.011 | ||
| Deviance | 1857.37 | 1921.71 | ||
| Null Deviance | 1908.84 | 1942.18 | ||
| Log likelihood | -928.68 | -960.86 | ||
| Akaike Inf. Crit. | 1871.37 | 1935.71 | ||